OpenStudy (jessicawade):

1 Trigonometry question, help please?

OpenStudy (jessicawade):

The coordinates of pi/3 on the unit circle are (1/2, sqr3/2) For an angle in Quadrant IV in the coordinate plane that has the same reference angle as pi/3, what is the positive radian measure of the angle and the corresponding coordinates?

OpenStudy (jessicawade):

@ganeshie8 sorry for the tag, do you know how to do this?

ganeshie8 (ganeshie8):

simply subtract pi/3 from 2pi : \[2\pi - \frac{\pi}{3}=?\]

OpenStudy (jessicawade):

oh lol

OpenStudy (jessicawade):

I got 5.23599 because I used a caculator

ganeshie8 (ganeshie8):

thats right, the exact value can be obtained by subtracting the fractions using a common denominator : \[2\pi - \frac{\pi}{3}=\frac{2\pi*3}{3}-\frac{\pi}{3}=\frac{6\pi-\pi}{3}=\frac{5\pi}{3}\]

ganeshie8 (ganeshie8):

thats a positive radian measure corresponding to the reference angle of pi/3 in fourth quadrant

OpenStudy (jessicawade):

so is it the 5.23599? or do I need to solve the equation?